# Assessing the dynamical assumptions in Tsirelson inequality tests of non-classicality in harmonic oscillators

### v2

@@ I.1 Outline and Preliminaries
This appendix is devoted to proving that if Tsirelson’s inequality is violated by a given quantum state, with support entirely in $\mathcal{H}_{k}$, then that state will satisfy all four LG3s (Eqs 1.12-1.15) if the measurement times $t_{1},t_{2},t_{3}$ are the same as in the Tsirelson scheme.
@@ I.2 Tsirelson Violation $\implies L_{1}>0$
$\langle Q_{1}+Q_{2}+Q_{3}\rangle>1$ (I1)
or in fact $\langle Q_{1}+Q_{2}+Q_{3}\rangle<-1$ (this doesn’t change the argument at all). Now, consider the Cauchy-Schwarz inequality applied to the left hand side of Equation I1
@@ I.3 $L_{1}$ is the smallest LG3 on $\mathcal{H}_{0},\mathcal{H}_{1}$ and $\mathcal{H}_{2}$
$C_{12}=\frac{1}{2}\langle\psi_{k}|Q(\omega^{-k}U+\omega^{k}U^{\dagger})Q|\psi_{k}\rangle\ ,$ (I5)
where $Q|\psi_{k}\rangle$ can be decomposed into $Q|\psi_{k}\rangle=|\phi_{k}\rangle+|\phi_{\perp}\rangle$, $|\phi_{k}\rangle\in\mathcal{H}_{k}$ and $\langle\phi_{k}|\phi_{\perp}\rangle=0$. Hence, as $|\phi_{\perp}\rangle$ is constructed out of orthogonal energy eigenstates to $|\phi_{k}\rangle$, we can write
@@
Finally, using the upper bound on $A||$ from Ref. [9],
@@
$\langle A\rangle\equiv\langle\hat{Q}_{1}+\hat{Q}_{2}+\hat{Q}_{3}\rangle$ (J1)
@@
Appendix K Probability Currents
K.1 Relating the Tsirelson quantity to the quantum current
Consider the limit definition of the integral to write $\langle\hat{T}_{D}\rangle$ as a Riemann sum,
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{2\pi}\left[\lim_{N\rightarrow\infty}\frac{2\pi}{N+1}\sum_{n=0}^{N}\left\langle\theta\left(\frac{2\pi n}{N+1}\right)\right\rangle\right]\ ,$ (K1)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{3}\right)\right\rangle+\frac{\varepsilon}{2\pi}\ ,$ (K2)
@@
Assuming $\langle\theta(t)\rangle$ is analytic in $t$, then $\langle\theta(t)\rangle$ has a well-defined Taylor expansion and we can borrow some well-known results from numerical analysis to quantify our error term. Specifically, we notice that this amounts to a left endpoint approximation to the integral for the dwell time. Therefore we can use the formula for the error term [76],
$E_{L}=\frac{(2\pi)^{2}}{6}\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau},\tau\in[0,2\pi]\ ,$ (K3)
@@
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{3}\right)\right\rangle+\frac{E_{L}}{2\pi}$ (K4)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(\frac{2\pi n}{3}\right)\right\rangle+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}\ ,$ (K5)
where $\tau\in[0,2\pi]$. The sum can be related to the Tsirelson quantity because $\left\langle\theta\right\rangle=\frac{\left\langle Q\right\rangle+1}{2}$ and $\langle A\rangle=\sum_{n=0}^{2}\left\langle Q\right\rangle$. So,
$\langle\hat{T}_{D}\rangle$ $=\frac{1}{6}\left(\sum_{n=0}^{2}\left\langle Q\left(\frac{2\pi n}{3}\right)\right\rangle+3\right)+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}$ (K6)
$=\frac{1}{6}\left(3+\langle A\rangle\right)+\frac{2\pi}{6}\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}.$ (K7)
@@
$\langle A\rangle=(6\langle\hat{T}_{D}\rangle-3)-2\pi\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}=-2\pi\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}$ (K8)
@@
Note that because $\langle\theta(t+\pi)\rangle=1-\langle\theta(t)\rangle$, the derivatives of $\langle\theta(t)\rangle$ simply flip sign after half a period. Hence, without loss of generality, we may drop the negative signs on our expressions, since $\tau$ is some unknown time on the interval $0,2\pi]$:
$\langle A\rangle=2\pi\frac{d\langle\theta\rangle}{dt}\big|_{t=\tau}$}\ . (K9)
We can rewrite Eq. K9 as,
$\langle A\rangle=2\pi\frac{d}{dt}\int_{0}^{\infty}|\psi(x,\tau)|^{2}dx.$ (K10)
@@
$\frac{\partial\rho}{\partial t}+\frac{\partial J}{\partial x}=0,$ (K11)
where $\rho=|\psi|^{2}$ is the probability density and $J=\operatorname{Im}\left(\psi^{*}\partial_{x}\psi\right)$ is the probability density current [77], we can substitute the current into the Tsirelson quantity:
$\langle A\rangle$ $=-2\pi\int_{0}^{\infty}\frac{\partial J(x,\tau)}{\partial x}dx,$ (K12)
$=2\pi J(0,\tau).$ (K13)
@@
$\exists\tau\in[0,2\pi]\text{ }\operatorname{s.t}\text{ }\langle A\rangle=2\pi J(0,\tau).$ (K14)
From Equation K14, we can make the following claim: if $|J(0,\tau)|\leq\frac{1}{2\pi},\forall\tau\in[0,2\pi]$, then $|\langle A\rangle|\leq 1$. That is to say, if the current at the origin stays sufficiently low, then we can guarantee the state will not violate the Tsirelson inequality. Notably, $J(0,t)$ could be measured by taking the time derivative of measured values of $\langle\theta(t)\rangle$ which only requires single-time measurements.
@@
K.2 A Tsirelson satisfaction condition on $\mathcal{H}_{0}$
@@
$T_{2/3}$ $=\frac{1}{6}\left(\sum_{j=1}^{3}\langle\theta_{j}\rangle+\langle\theta_{1}\rangle\right)+\frac{1}{2\pi}\sum_{j=1}^{3}E_{j}$ (K16)
$=\frac{1}{12}\langle A\rangle+\frac{1}{12}\langle Q_{1}\rangle+\frac{1}{3}+\frac{1}{2\pi}\sum_{j=1}^{3}E_{j}.$ (K17)
@@
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-\frac{6}{\pi}\sum_{j=1}^{3}E_{j}$ (K18)
where both $E_{1}$ and $E_{2}$ are left end point errors with $E_{1}=\frac{(2\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta\rangle}{\mathrm{d}t}\big|_{\tau_{1}}\ ,\ \tau_{1}\in[0,2\pi/3]$ and $E_{2}=\frac{(\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta\rangle}{\mathrm{d}t}\big|_{\tau_{2}}\ ,\ \tau_{2}\in[2\pi/3,\pi]$. However, $E_{3}$ is a right endpoint error with $E_{3}=-\frac{(\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta\rangle}{\mathrm{d}t}\big|_{\tau_{3}}\ ,\ \tau_{3}\in[\pi,4\pi/3]$. So, expanding out and simplifying the error term contribution,
$\sum_{j=1}^{3}E_{j}=\frac{\pi^{2}}{18}\left[4J(\tau_{1})+J(\tau_{2})-J(\tau_{3})\right]\ .$ (K19)
@@
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-\frac{\pi}{3}\left[4J(\tau_{1})+J(\tau_{2})-J(\tau_{3})\right]\ .$ (K20)
But, by noting that $J(\tau_{3})=+J(\tilde{\tau}_{3}),\ \tilde{\tau}_{3}\in[0,\pi/3]$, the sum of the three currents can be reformulated as a weighted average over the domain $0,\pi]$. Pairing that with an application of the mean value theorem allows us to write
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-2\pi J(\tau),\ \tau\in[0,\pi]\ .$ (K21)
@@
$\langle A\rangle=-\frac{3\pi}{2}J(\tau),\ \tau\in[0,2\pi/3]\ ,$ (K22)
@@
$\text{For}\ |\psi\rangle\in\mathcal{H}_{0}\ ,\ \exists\tau\in[0,2\pi/3]\ \text{s.t}\ \langle A\rangle=\frac{3\pi}{2}J(\tau).$ (K23)
@@
$\text{For}\ |\psi\rangle\in\mathcal{H}_{0}\ ,\ \text{if}\ |J(\tau)|\leq\frac{2}{3\pi}\ \forall\tau\in[0,2\pi/3]\ \text{then}\ |\langle A\rangle|\leq 1\ .$ (K24)
Curiously, when we computed the classical probability current, using
$j_{c}(0,t)=\iint\mathrm{d}x\mathrm{d}p\ p(t)\delta[x(t)]W_{c}(x,p)\ ,$ (K25)
for any natural classicalisation of the phase-space distribution, $W_{c}(x,p)$, associated with $|\psi_{Z}\rangle\in\mathcal{H}_{0}$, we found the classical current was lower in magnitude than the quantum current across the entire period. Specifically, from the Wigner function of the Zaw-Scarani state, $W_{Z}(x,p)$, we constructed the following classical distributions: (1) $W_{c}(x,p)\propto|\langle\alpha|\psi_{Z}\rangle|^{2}$ (the Husimi-Q distribution), (2) $W_{c}(x,p)\propto|W_{Z}(x,p)|$, (3) $W_{c}(x,p)\propto W_{Z}(x,p)+|W_{Z}(x,p)|$ and (4) $W_{c}(x,p)=|\psi_{Z}(x)|^{2}|\phi_{Z}(p)|^{2}$ (where $\phi_{Z}(p)$ is the momentum-space wavefunction of the Zaw-Scarani state). In Figure 9, we illustrate the resulting classical current, where the classicalisation strategy chosen is using the Husimi-Q distribution, and compare it to the quantum current. This result perhaps suggests a connection between the current based bound on the Tsirelson quantity in Eq. K24 and classicalised currents of violating quantum states.

### v3

@@ I.1 Outline and Preliminaries
This appendix is devoted to proving that if Tsirelson’s inequality is violated by a given quantum state, with support entirely in $\mathcal{H}_{k}$, then that state will satisfy all four LG3s (Eqs 2.9-2.12) if the measurement times $t_{1},t_{2},t_{3}$ are the same as in the Tsirelson scheme.
@@ I.2 Tsirelson Violation $\implies L_{1}>0$
$|\langle Q_{1}+Q_{2}+Q_{3}\rangle|>1$ (I1)
Now, consider the Cauchy-Schwarz inequality applied to the left hand side of Eq. I1
@@ I.3 $L_{1}$ is the smallest LG3 on $\mathcal{H}_{0},\mathcal{H}_{1}$ and $\mathcal{H}_{2}$
$C_{12}=\frac{1}{2}\langle\psi_{k}|\hat{Q}(\omega^{-k}U+\omega^{k}U^{\dagger})\hat{Q}|\psi_{k}\rangle\ ,$ (I5)
where $\hat{Q}|\psi_{k}\rangle$ can be decomposed into $\hat{Q}|\psi_{k}\rangle=|\phi_{k}\rangle+|\phi_{\perp}\rangle$, $|\phi_{k}\rangle\in\mathcal{H}_{k}$ and $\langle\phi_{k}|\phi_{\perp}\rangle=0$. Hence, as $|\phi_{\perp}\rangle$ is constructed out of orthogonal energy eigenstates to $|\phi_{k}\rangle$, we can write
@@
Finally, using the upper bound on $|\hat{A}$$|$ from Ref. [zaw2024threeanglevariantstsirelsonsprecession],
@@
$\langle A\rangle\equiv\langle{Q}_{1}+{Q}_{2}+{Q}_{3}\rangle$ (J1)
@@
Appendix K Smoothed Projectors
[…]
The smoothing will also affect the probabilities $p(+,+,+)$ and p(,,)p(-,-,-) characterizing uniform precession. To assess this we use a simplified but instructive “smoothing” function $\theta_{\epsilon}(x)$, which agrees with the usual $\theta(x)$ function except that it takes the value $\frac{1}{2}$ for $\epsilon/2<x<\epsilon/2$. Also, for simplicity, we use a classical analysis with a state concentrated around $x=0$ and $p=0$.
Then there are two cases. If the width of the state is less than $\epsilon$, then most of the state is concentrated in the smoothing region and we have effectively $\theta_{\epsilon}(x)=\frac{1}{2}$. We would then have $p(+,+,+)$ be of order $1/8$, which means that uniform precession is not satisfied well. However, if the width of the state is much greater $\epsilon$, then only a small fraction of the state will be in the smoothing region and the above effect will be proportionately reduced. For the Zaw–Scarani state, plotted in Fig. 9, the proportion of the state found within distance $\delta$ of the origin is plotted in Fig. 9. We see that this is very small for $\delta\ll 1$. This means that uniform precession holds up well for smoothing with $\epsilon\ll 1$, the same region in which the Tsirelson violation is maintained. These order of magnitude arguments clearly need to be confirmed by a more detailed calculation, similar to that of Ref. [MawbySHO], but we will not do this here.
Figure 9: (a) A plot of $|\psi(x)|^{2}$ against the dimensionless position for the Zaw-Scarani state $|\Psi_{\mathcal{Z}}\rangle=\frac{4}{\sqrt{42}}|0\rangle-\frac{1}{\sqrt{2}}|3\rangle+\sqrt{\frac{5}{42}}|6\rangle$ (b) For the Zaw-Scarani state, this plot shows the probability of measuring the particle to be in the interval $(-\delta,\delta)$ against $\delta$, where $\delta$ is the dimensionless position.
Appendix L Probability Currents
L.1 Relating the Tsirelson quantity to the quantum current
Consider the limit definition of the integral to write $\langle T_{D}(2\pi)\rangle$ as a Riemann sum,
$\langle T_{D}(2\pi)\rangle$ $=\frac{1}{2\pi}\left[\lim_{N\rightarrow\infty}\frac{2\pi}{N+1}\sum_{n=0}^{N}\left\langle\theta\left(x\left(\frac{2\pi n}{N+1}\right)\right)\right\rangle\right]\ ,$ (L1)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(x\left(\frac{2\pi n}{3}\right)\right)\right\rangle+\frac{\varepsilon}{2\pi}\ ,$ (L2)
@@
Assuming $\langle\theta(x(t))\rangle$ is continuously differentiable in $t$, then Taylor’s theorem applies to $\langle\theta(x(t))\rangle$ and we can borrow some well-known results from numerical analysis to quantify our error term. Specifically, we notice that this amounts to a left endpoint approximation to the integral for the dwell time. Therefore we can use the formula for the error term [numerical_methods],
$E_{L}=\frac{(2\pi)^{2}}{6}\frac{d\langle\theta(x)\rangle}{dt}\big|_{t=\tau},\tau\in[0,2\pi]\ ,$ (L3)
@@
$\langle T_{D}(2\pi)\rangle$ $=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(x\left(\frac{2\pi n}{3}\right)\right)\right\rangle+\frac{E_{L}}{2\pi}$ (L4)
$=\frac{1}{3}\sum_{n=0}^{2}\left\langle\theta\left(x\left(\frac{2\pi n}{3}\right)\right)\right\rangle+\frac{2\pi}{6}\frac{d\langle\theta(x)\rangle}{dt}\big|_{t=\tau}\ ,$ (L5)
where $\tau\in[0,2\pi]$. The sum can be related to the Tsirelson quantity because $\left\langle\theta(x(t))\right\rangle=\frac{\left\langle Q(t)\right\rangle+1}{2}$ and $\langle A\rangle=\sum_{n=0}^{2}\left\langle Q\right\rangle$. So,
$\langle T_{D}(2\pi)\rangle$ $=\frac{1}{6}\left(\sum_{n=0}^{2}\left\langle Q\left(\frac{2\pi n}{3}\right)\right\rangle+3\right)+\frac{2\pi}{6}\frac{d\langle\theta(x)\rangle}{dt}\big|_{t=\tau}$ (L6)
$=\frac{1}{6}\left(3+\langle A\rangle\right)+\frac{2\pi}{6}\frac{d\langle\theta(x)\rangle}{dt}\big|_{t=\tau}.$ (L7)
@@
$\langle A\rangle=(6\langle T_{D}(2\pi)\rangle-3)-2\pi\frac{d\langle\theta(x)\rangle}{dt}\big|_{t=\tau}=-2\pi\frac{d\langle\theta(x)\rangle}{dt}\big|_{t=\tau}$ (L8)
@@
Note that because $\langle\theta(x(t+\pi))\rangle=1-\langle\theta(x(t))\rangle$, the derivatives of $\langle\theta(x(t))\rangle$ simply flip sign after half a period. Hence, without loss of generality, we may drop the negative signs on our expressions, since $\tau$ is some unknown time on the interval $0,2\pi]$:
$\langle A\rangle=2\pi\frac{d\langle\theta(x)\rangle}{dt}\big|_{t=\tau}$}\ . (L9)
We can rewrite Eq. L9 as,
$\langle A\rangle=2\pi\frac{d}{dt}\int_{0}^{\infty}|\psi(x,\tau)|^{2}dx.$ (L10)
@@
$\frac{\partial\rho}{\partial t}+\frac{\partial J}{\partial x}=0,$ (L11)
where $\rho=|\psi|^{2}$ is the probability density and $J=\operatorname{Im}\left(\psi^{*}\partial_{x}\psi\right)$ is the probability density current [Sakurai_Napolitano_2020], we can substitute the current into the Tsirelson quantity:
$\langle A\rangle$ $=-2\pi\int_{0}^{\infty}\frac{\partial J(x,\tau)}{\partial x}dx,$ (L12)
$=2\pi J(0,\tau).$ (L13)
@@
$\exists\tau\in[0,2\pi]\text{ }\operatorname{s.t}\text{ }\langle A\rangle=2\pi J(0,\tau).$ (L14)
From Eq. L14, we can make the following claim: if $|J(0,\tau)|\leq\frac{1}{2\pi},\forall\tau\in[0,2\pi]$, then $|\langle A\rangle|\leq 1$. That is to say, if the current at the origin stays sufficiently low, then we can guarantee the state will not violate the Tsirelson inequality. Notably, $J(0,t)$ could be measured by taking the time derivative of measured values of $\langle\theta(x(t))\rangle$ which only requires single-time measurements.
@@
L.2 A Tsirelson satisfaction condition on $\mathcal{H}_{0}$
@@
$T_{2/3}$ $=\frac{1}{6}\left(\sum_{j=1}^{3}\langle\theta(x(t_{j}))\rangle+\langle\theta(x(t_{1}))\rangle\right)+\frac{1}{2\pi}\sum_{j=1}^{3}E_{j}$ (L16)
$=\frac{1}{12}\langle A\rangle+\frac{1}{12}\langle Q_{1}\rangle+\frac{1}{3}+\frac{1}{2\pi}\sum_{j=1}^{3}E_{j}.$ (L17)
@@
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-\frac{6}{\pi}\sum_{j=1}^{3}E_{j}$ (L18)
where both $E_{1}$ and $E_{2}$ are left end point errors with $E_{1}=\frac{(2\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta(x)\rangle}{\mathrm{d}t}\big|_{\tau_{1}}\ ,\ \tau_{1}\in[0,2\pi/3]$ and $E_{2}=\frac{(\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta(x)\rangle}{\mathrm{d}t}\big|_{\tau_{2}}\ ,\ \tau_{2}\in[2\pi/3,\pi]$. However, $E_{3}$ is a right endpoint error with $E_{3}=-\frac{(\pi/3)^{2}}{2}\frac{\mathrm{d}\langle\theta(x)\rangle}{\mathrm{d}t}\big|_{\tau_{3}}\ ,\ \tau_{3}\in[\pi,4\pi/3]$. So, expanding out and simplifying the error term contribution,
$\sum_{j=1}^{3}E_{j}=\frac{\pi^{2}}{18}\left[4J(\tau_{1})+J(\tau_{2})-J(\tau_{3})\right]\ .$ (L19)
@@
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-\frac{\pi}{3}\left[4J(\tau_{1})+J(\tau_{2})-J(\tau_{3})\right]\ .$ (L20)
But, by noting that $J(\tau_{3})=+J(\tilde{\tau}_{3}),\ \tilde{\tau}_{3}\in[0,\pi/3]$, the sum of the three currents can be reformulated as a weighted average over the domain $0,\pi]$. Pairing that with an application of the intermediate value theorem allows us to write
$\langle A\rangle=12T_{2/3}-4-\langle Q_{1}\rangle-2\pi J(\tau),\ \tau\in[0,\pi]\ .$ (L21)
@@
$\langle A\rangle=-\frac{3\pi}{2}J(\tau),\ \tau\in[0,2\pi/3]\ ,$ (L22)
@@
$\text{For}\ |\psi\rangle\in\mathcal{H}_{0}\ ,\ \exists\tau\in[0,2\pi/3]\ \text{s.t}\ \langle A\rangle=\frac{3\pi}{2}J(\tau).$ (L23)
@@
$\text{For}\ |\psi\rangle\in\mathcal{H}_{0}\ ,\ \text{if}\ |J(\tau)|\leq\frac{2}{3\pi}\ \forall\tau\in[0,2\pi/3]\ \text{then}\ |\langle A\rangle|\leq 1\ .$ (L24)
L.3 Classical and Quantum Probability Current
Peculiarly, when we computed the classical probability current, using
$j_{c}(0,t)=\iint\mathrm{d}x\mathrm{d}p\ p(t)\delta[x(t)]W_{c}(x,p)\ ,$ (L25)
for any natural classicalisation of the phase-space distribution, $W_{c}(x,p)$, associated with $|\psi_{Z}\rangle\in\mathcal{H}_{0}$, we found the classical current was lower in magnitude than the quantum current across the entire period. Specifically, from the Wigner function of the Zaw-Scarani state, $W_{Z}(x,p)$, we constructed the following classical distributions: (1) $W_{c}(x,p)\propto|\langle\alpha|\psi_{Z}\rangle|^{2}$ (the Husimi-Q distribution), (2) $W_{c}(x,p)\propto|W_{Z}(x,p)|$, (3) $W_{c}(x,p)\propto W_{Z}(x,p)+|W_{Z}(x,p)|$ and (4) $W_{c}(x,p)=|\psi_{Z}(x)|^{2}|\phi_{Z}(p)|^{2}$ (where $\phi_{Z}(p)$ is the momentum-space wavefunction of the Zaw-Scarani state). In Figure 10, we illustrate the resulting classical current, where the classicalisation strategy chosen is using the Husimi-Q distribution, and compare it to the quantum current. This result perhaps suggests a connection between the current based bound on the Tsirelson quantity in Eq. L24 and classicalised currents of violating quantum states.